OpenStudy (anonymous):

Calculate x-60/x = 0.40

OpenStudy (anonymous):

blexting plz help me slove my qestion

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

x - 60 40 ----- = --- x 100 anytime you have a fraction = fraction you cross multiply and then solve. 100(x - 60) = 40x distribute 100x - 6000 = 40x subtract 100x -6000 = -60x divide by -60 100 = x

OpenStudy (anonymous):

you just said subistud 4 for x

OpenStudy (anonymous):

Thank you blexting and mary1000.